# Freezing of Gait Detection Using Gramian Angular Fields and Federated Learning from Wearable Sensors

### v1

@@ 3.6. Experimental Setup
The first block utilizes 32 filters with 20% dropout, while the second and third blocks employ 64 and 128 filters respectively, incorporating L2 regularization ($\lambda=0.001$) and increasing dropout rates (20% and 40%). Each convolutional block integrates batch normalization and max pooling operations to enhance feature learning and reduce spatial dimensions. A global average pooling layer follows the convolutional blocks, effectively reducing parameters while capturing essential global features from each branch.
Features extracted from individual branches are concatenated to create a unified representation, which then passes through two dense layers for final classification. The first dense layer comprises 128 neurons with ReLU activation, L2 regularization ($\lambda=0.001$), and 40% dropout, followed by a second dense layer of 64 neurons with ReLU activation and 20% dropout. The final classification is performed by a softmax layer with two neurons corresponding to FOG and no-FOG classes. To ensure robust performance and prevent overfitting, the model incorporates multiple regularization techniques Adam optimizer with the categorical cross-entropy loss function. The backbone architecture remained unchanged during channel configuration analysis, ensuring performance differences stemmed solely from input channel variations.
@@
FOGSense was evaluated using standard binary classification metrics. Based on the confusion matrix elements—true positives (TP), false positives (FP), false negatives (FN), and true negatives (TN)—we computed the following metrics: sensitivity, precision, accuracy, and F1-score.
@@
4. Results
Our experiments excluded conventional data augmentation to preserve GAF signal integrity. We evaluated window-level and temporal FOG detection, achieving promising results. To enhance clinical applicability, we used federated learning for continuous model adaptation, preserving privacy and enabling model refinement across edge devices.
Table 2. FOGSense’s performance for three different detection types using all three channels. DFW: Detected FOG Windows, DFE: Detected FOG Episodes.
Detection Type Configuration Accuracy (%) F1 (%)
Window-level (DFW) All Channels 87.23 90.31
Episode-level (DFE) All Channels 87.23 90.86
Federated Learning All Channels 86.98 90.47
4.1. Effect of Channel Selection
@@
Table 3. Impact of selection of AccV (Vertical), AccML (Mediolateral), and AccAP (Anteroposterior) axes.
@@
Table 3 reveals several key insights regarding channel configuration impact on model performance. The baseline three-channel configuration (AccAP-AccML-AccV) achieved 87.23% accuracy and 90.31% F1-score, requiring 63 seconds per training epoch. However, reducing to two channels generally improved performance while halving the training time to approximately 32 seconds per epoch. Among dual-channel configurations, AccML-AccV demonstrated superior performance with 95.30% accuracy and 90.75% F1-score.
@@
Figure 7. Comparison of channel configurations in the FOGSense system shows that the vertical axis (AccV) obtains a better F1 score than other channel choices.
@@
4.2. Comparison with State-of-the-Art
@@
We present a comparative analysis of FOGSense and the other state-of-the-art (SOTA) methods in Table 4. Our FOGSense method achieves an F1 score of 0.96 on the ‘tdcsfog’ dataset, whereas the closest competitor LIFT-PD achieved an F1 score of 0.79. Moreover, unlike the other methods, the FOGSense maintains a score of at least 0.90 in all four metrics. Additionally, FOGSense’s use of the vertical axis (AccV) channel selection not only enhances detection accuracy but also provides a more efficient approach to gait monitoring. These results underscore FOGSense’s potential as a powerful and dependable tool for real-time gait analysis.
@@
Table 4. FOGSense vs. state-of-the-art on tdcsfog. The best values for each metric are highlighted in bold, while the second-best values are underlined.
Study DFE SEN (DFW) PRE F1
One Class Classifier (Naghavi and Wade, 2021) 90.8% 71.6% 0.86 0.77
Semi-Supervised Model (Mikos et al., 2017) 72.3%
Multi-head CNN (Borzì et al., 2023) 97.3% 94.6% 0.56 0.68
LIFT-PD 88% 84% 0.74 0.79
FOGSense 96.3% 0.958 0.963
5. Discussion
@@
In this paper, we introduced a novel approach to detect FOG by integrating GAF-transformed accelerometer data with a multichannel CNN architecture and a decentralized federated learning pipeline. Our experiments demonstrated not only high detection accuracy but also highlighted the practicality of our system for real-world deployment. A key finding is that a single-channel configuration using only the vertical acceleration component (AccV) achieved impressive performance—96.25% accuracy and 96.27% F1-score—while significantly reducing computational demands compared to multi-channel configurations. This result challenges the traditional view that multiple sensor channels are essential for accurate FOG detection.
@@
To address continuous monitoring constraints, we developed a distributed framework that synchronizes data between wearable devices and smartphones during off-peak hours, balancing the demands of monitoring, battery life, and computational efficiency. Although the GAF transformation was effective in our approach, exploring alternative time series representations such as Recurrence Plots and Markov Transition Fields may yield further insights. Additionally, while our distributed framework shows promise in simulations, real-world validation on multiple devices is a critical next step. Investigating different weight calculation methods beyond weighted averaging could enhance the learning dynamics in federated learning. Testing the model across diverse patient populations and datasets would further strengthen its generalizability, and studying patient data mixing strategies may provide insights into managing individual gait variations effectively.
6. Conclusion
@@
Accurate detection of freezing of gait (FOG) can significantly enhance early intervention strategies, improve mobility support, and ultimately contribute to a better quality of life for individuals with Parkinson’s disease. Our FOGSense system uses Gramian Angular Field (GAF) imaging, convolutional neural networks (CNN), and federated learning to ensure accurate and resource-efficient FOG detection in an uncontrolled environment. Moreover, FOGSense is robust to sensor failure as it needs only one channel to make an accurate prediction and it can delegate the prediction task to a different single-channel model by dynamic weight transfer in case of a channel failure.
@@
Our work establishes a foundation for efficient, practical, and reliable FOG detection systems. We provide new insights into lightweight, single-channel solutions that outperform state-of-the-art multi-channel solutions. We also identified future research directions that could advance the field, from alternative data representations to federated learning enhancements and real-world validations. This study highlights the potential of FOGSense and opens avenues for refining FOG detection systems to meet clinical and practical demands more effectively.

### v3

@@ III-F Experimental Setup
Batch normalization and max pooling enhance feature extraction while reducing spatial dimensions.
A global average pooling layer further compresses parameters before features from all branches are concatenated. The final classifier consists of two dense layers (128 and 64 neurons with ReLU activation, dropout (40%, 20%), and L2 regularization) followed by a softmax layer for FOG classification. Adam optimizer and categorical cross-entropy loss are used to improve generalization. The backbone architecture remained unchanged during channel configuration analysis to ensure consistent performance evaluation.
@@
III-G Performance Measures
FOGSense was evaluated using standard binary classification metrics. Based on the confusion matrix elements—true positives (TP), false positives (FP), false negatives (FN), and true negatives (TN)—we computed the following metrics: sensitivity, precision, accuracy, F1-score, and false positive rate (FPR). FPR is crucial in FOG detection as it quantifies incorrect intervention triggers, ensuring the system minimizes unnecessary alerts; otherwise, treatment efficacy will decrease.
@@ III-G Performance Measures
IV Results
To ensure the integrity of GAF-transformed signals, we did not apply conventional data augmentation techniques. We evaluated FOGSense for both window-level and episode-level detection, achieving consistently high performance. Window-level detection refers to independently classifying each window as FOG or non-FOG and episode-level detection merges contiguous windows to form a single FOG event. To improve clinical applicability, we implemented federated learning, allowing continuous model adaptation while preserving privacy. Table II shows that FOGSense achieves high accuracy across all detection levels, with F1-scores above 90% and a low false positive rate (3.2-3.8%). Episode-level detection performs best, minimizing false alarms.
TABLE II: FOGSense’s performance for three different detection types using all three channels. DFW: Detected FOG Windows, DFE: Detected FOG Episodes, FPR: False Positive Rate.
Detection Type Configuration Accuracy (%) F1 (%) FPR (%)
Window-level (DFW) All Channels 87.23 90.31 3.8
Episode-level (DFE) All Channels 87.23 90.86 3.2
Federated Learning All Channels 86.98 90.47 3.5
IV-A Effect of Channel Selection
@@
TABLE III: Impact of channel selection
@@
Table III reveals several key insights regarding channel configuration impact on model performance. The baseline three-channel configuration (AccAP-AccML-AccV) achieved 87.23% accuracy and 90.31% F1-score, requiring 63 seconds per training epoch. However, reducing to two channels generally improved performance while halving the training time to approximately 32 seconds per epoch. Among dual-channel configurations, AccML-AccV demonstrated superior performance with 95.30% accuracy and 90.75% F1-score.
@@
Figure 7: Comparison of channel configurations in the FOGSense system shows that the vertical axis (AccV) obtains a better F1 score than other channel choices.
@@
IV-B Comparison with State-of-the-Art
@@
We present a comparative analysis of FOGSense and the other state-of-the-art (SOTA) methods in Table IV. Our FOGSense method achieves an F1 score of 0.96 on the ‘tdcsfog’ dataset, whereas the closest competitor, LIFT-PD achieved an F1 score of 0.79. Moreover, unlike the other methods, the FOGSense maintains a score of at least 0.90 in all four metrics, with an average 74.5% reduction in FPR. Additionally, FOGSense’s use of the vertical axis (AccV) channel selection not only enhances detection accuracy but also provides a more efficient approach to gait monitoring. These results underscore FOGSense’s potential as a powerful and dependable tool for real-time gait analysis.
@@
TABLE IV: FOGSense vs. state-of-the-art on tdcsfog. The best values are highlighted in bold, while the second-best values are underlined.
Study DFE SEN (DFW) PRE F1 FPR
One Class Classifier [27] 90.8% 71.6% 0.86 0.77 12.4%
Semi-Supervised Model [39] 72.3% 11.8%
Multi-head CNN [40] 97.3% 94.6% 0.56 0.68 20.5%
LIFT-PD [29] 88% 84% 0.74 0.79 9.6%
FOGSense 96.3% 0.958 0.963 3.2%
V Discussion
@@
In this paper, we introduced a novel approach to detect FOG by integrating GAF-transformed accelerometer data with a multichannel CNN architecture and a decentralized federated learning pipeline. Our experiments demonstrated not only high detection accuracy with a low false positive intervention rate but also highlighted the practicality of our system for in-home real-time monitoring. A key finding is that a single-channel configuration using only the vertical acceleration component (AccV) achieved impressive performance—96.25% accuracy and 96.27% F1-score—while significantly reducing computational demands compared to multi-channel configurations. This result challenges the traditional view that multiple sensor channels are essential for accurate FOG detection while ensuring a balance between high accuracy and minimal false positives, preventing unnecessary interventions that could reduce treatment efficacy over time.
@@
To address continuous monitoring constraints, we developed a distributed framework that synchronizes data between wearable devices and smartphones during off-peak hours, balancing the demands of monitoring, battery life, and computational efficiency. Although the GAF transformation was effective in our approach, exploring alternative time series representations such as Recurrence Plots and Markov Transition Fields may yield further insights. Additionally, while our distributed framework shows promise in simulations, real-world validation on multiple edge devices is a critical next step. Investigating different weight calculation methods beyond weighted averaging could enhance the learning dynamics in federated learning. Testing the model across diverse patient populations and datasets would further strengthen its generalizability, and studying patient data mixing strategies may provide insights into managing individual gait variations effectively.
VI Conclusion
@@
Accurate detection of freezing of gait (FOG) can significantly enhance early intervention strategies, improve mobility support, and ultimately contribute to a better quality of life for individuals with Parkinson’s disease. Our FOGSense system uses Gramian Angular Field (GAF) imaging, CNN, and federated learning to ensure accurate and resource-efficient FOG detection in an uncontrolled environment. Moreover, FOGSense maintains reliability even with sensor failure, as it needs only one channel to make an accurate prediction and it can delegate the prediction task to a different single-channel model by dynamic weight transfer in case of a channel failure.
@@
Our work establishes a foundation for efficient, practical, and reliable FOG detection systems. We provide new insights into lightweight, single-channel solutions that outperform multi-channel approaches, achieving a 22.2% improvement in F1-score and a 74.5% reduction in false positive rate (FPR), minimizing unnecessary interventions that could diminish their therapeutic impact over time. By reducing false positives, FOGSense ensures that patients receive interventions only when needed, preventing desensitization to stimuli and maintaining long-term clinical effectiveness. We also identified future research directions that could advance the field, from alternative data representations to federated learning enhancements and real-world validations. This study highlights the potential of FOGSense and opens avenues for refining real-time FOG detection systems to meet clinical and practical demands more effectively.